# Parametrically enhancing sensor sensitivity at an exceptional point

### 12 Dec 2023 v2

@@ Appendix B Splitting vs shifting sensing schemes
This Appendix is meant to provide a comparison between the two sensing schemes investigated in the main text. We have shown the possibility of sensing based either the splitting or the shifting of the $\rm{EP}$, but a straight comparison between the two is missing. A general eigenvalue difference evaluated at the $\rm{EP}$ can be given by,
$\Delta\lambda_{\pm}^{\rm{EP}}=\frac{\Delta\omega^{\epsilon}}{2}-\frac{i|\delta\gamma|}{4}\pm\frac{\sigma_{\rm{EP}}^{\epsilon}}{2},$ (44)
@@
$\Delta\omega^{\epsilon}=\frac{1}{2}\left[\frac{\tilde{\omega}_{2}\delta\omega_{2}}{\Delta_{m}^{2}}-\frac{\chi_{2}\delta\chi_{2}}{\Delta_{m}^{2}}-\frac{\chi_{1}\delta\chi_{1}}{\Delta_{m}^{1}}-\frac{1}{2}\left(\frac{\chi_{1}^{2}}{\tilde{\omega}_{1}}+\frac{\chi_{2}^{2}}{\tilde{\omega}_{2}}\right)\right],$ (45)
@@
$\sigma_{\rm{EP}}^{\epsilon}=\sqrt{(8J_{m}r)^{2}+4\nu^{2}-4i\nu|\delta\gamma|+4\nu\mu-2i\mu|\delta\gamma|},$ (46)
with $\nu=\frac{1}{2}\left(\frac{\chi_{2}^{2}}{\tilde{\omega}_{2}}-\frac{\chi_{1}^{2}}{\tilde{\omega}_{1}}\right)+\frac{\chi_{2}\delta\chi_{2}}{\Delta_{m}^{2}}-\frac{\tilde{\omega}_{2}\delta\omega_{2}}{\Delta_{m}^{2}}-\frac{\chi_{1}\delta\chi_{1}}{\Delta_{m}^{1}}$ and $\mu=(2\delta\tilde{\omega}-i\delta\Gamma_{\rm{eff}})$. In these expressions, $\delta\omega_{2}$ and $\delta\gamma$ are the two types of perturbation used in the main text. The former is induced by a mass deposition on the second mechanical resonator, while the latter comes from optical fluctuations or a mismatch damping of the two mechanical resonators. We stress that there can be also a perturbation induced by the fluctuation of the parametric driving which can be captured through $\delta\chi_{j}$ in Eq. 46. However, we have not consider such a perturbation in our investigation in the main text. We have also made some assumptions for a seek of simplicity, $\tilde{\omega}_{1}=\tilde{\omega}_{2}=\omega_{m}$ and $\chi_{1}=\chi_{2}=\chi$. With these approximations, we were able to derive the different expressions for sensitivity and enhancement factor presented in the main text, which were qualitatively in good agreement with our results. In order to give a direct comparison between the two sensing scenarios, we plotted both their sensitivity and enhancement factor in Fig. 7. The full line is from the sensing based on the shifting mechanism while the dashed line refers to splitting at the $\rm{EP}$. It can be clearly seen that the scheme based on the splitting performs better as discussed in the main text.
Figure 7: Comparison of the sensitivities and the enhancement factors for the two sensing schemes when $\chi=0$. The used parameters are those as in Fig. 7.
@@ Appendix C Transmission and output spectra
$\mathcal{O}(\omega)=\mathcal{D}(\omega)\left[i\sqrt{\mathcal{K}}\mathcal{N}(\omega)+\sqrt{\gamma_{m}}\mathcal{O}^{in}(\omega)\right],$ (47)
where $\mathcal{D}(\omega)=(-\mathcal{M}-i\omega)^{-1}$. We are interested to obtain the spectra two optical modes at the output of the opto-mechanical cavity. According to the input-output theory, the operators for the output fields are related to the cavity and to the input noise operators by the relation $\delta\alpha^{out}_{1}(\omega)=\sqrt{\kappa}\delta\alpha_{1}(\omega)-\delta\alpha^{in}_{1}(\omega)$ and $\delta\alpha^{out}_{2}(\omega)=\sqrt{\kappa}\delta\alpha_{2}(\omega)-\delta\alpha^{in}_{2}(\omega)$. Then by using Eq.(A) and Eq.(47), the expressions for the two transmitted fields in Fourier domain are given by,
$\delta\alpha^{out}_{1}(\omega)$ $=S_{11}(\omega)\delta\alpha^{in}_{1}(\omega)+S_{12}(\omega)\delta\alpha^{in^{\dagger}}_{2}(\omega)+S_{13}(\omega)\delta\beta^{in}_{1}(\omega)$
$+S_{14}(\omega)\beta^{in}_{2}(\omega),$ (48a)
$\alpha^{out^{\dagger}}_{2}(\omega)$ $=S_{21}(\omega)\delta\alpha^{in}_{1}(\omega)+S_{22}(\omega)\delta\alpha^{in^{\dagger}}_{2}(\omega)+S_{23}(\omega)\delta\beta^{in}_{1}(\omega)$
$+S_{24}(\omega)\beta^{in}_{2}(\omega)$ (48b)
@@
$S_{11}(\omega)=1-\dfrac{2\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{12}(\omega)=-\dfrac{2i\tilde{G}_{1}\tilde{J}_{m}\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (49a)
$S_{13}(\omega)=\dfrac{2i\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{14}(\omega)=\dfrac{2\tilde{G}_{1}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (49b)
$S_{21}(\omega)=\dfrac{2i\tilde{G}_{2}\tilde{J}_{m}\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{22}(\omega)=1+\dfrac{2\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{eff}(\omega)},$ (49c)
$S_{23}(\omega)=\dfrac{2\tilde{G}_{2}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)},\,\,S_{24}(\omega)=-\dfrac{2i\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{eff}(\omega)}$ (49d)
where $\xi_{eff}(\omega)=\xi_{1}(\omega)\xi_{2}(\omega)+\tilde{J}^{2}_{m}$ with $\xi_{1}(\omega)=\Gamma^{1}_{eff}/2+i(\Delta^{1}_{m}-\omega)$ and $\xi_{2}(\omega)=\Gamma^{2}_{eff}/2+i(\Delta^{2}_{m}-\omega)$. The input quantum noise operators cavity and mechanical modes are characterized by the following correlations
$\langle\delta\alpha^{in}_{j}(t),\delta\alpha^{in^{\dagger}}_{j}(t^{\prime})\rangle$ $=\delta(t+t^{\prime}),$ (50a)
$\langle\delta\beta^{in^{\dagger}}_{j}(t^{\prime}),\delta\beta^{in}_{j}(t)\rangle$ $=N_{s}\delta(t+t^{\prime}),$ (50b)
[…]
where $N^{j}_{s}=(n^{j}_{m}+1)\sinh^{2}(r_{j})+n^{j}_{m}\cosh^{2}(r)$ and $M^{j}_{s}=(2n^{j}_{m}+1)\sinh(r_{j})\cosh(r_{j})$ with $n^{j}_{m}=[e^{\hbar\omega_{j}/K_{B}T}-1]^{-1}$ is the mean thermal occupation number of the mechanical modes. Here we assumed that both mechanical resonators are at same temperature $T$. By using the expressions from Appendix C and Appendix C, one can derive the output spectra defined as
@@
$+\left(2N_{s_{2}}+1\right)\left|S_{24}(\omega)\right|{}^{2}.$ (51b)
where $N_{s_{1,2}}=(n_{m_{1,2}}+1)\sinh^{2}(r_{1,2})+n_{m_{1,2}}\cosh^{2}(r_{1,2})$.

### 13 Jan 2024 v3

@@ Appendix B Splitting vs shifting sensing schemes
This Appendix is meant to provide a comparison between the two sensing schemes investigated in the main text. We have shown the possibility of sensing based either the splitting or the shifting of the $\rm{EP}$, but a straight comparison between the two is missing. A general eigenvalue difference evaluated at the $\rm{EP}$ can be given by
$\Delta\lambda_{\pm}^{\rm{EP}}=\frac{\Delta\omega^{\epsilon}}{2}-\frac{{\rm i}|\delta\gamma|}{4}\pm\frac{\sigma_{\rm{EP}}^{\epsilon}}{2},$ (40)
@@
$\Delta\omega^{\epsilon}:=\frac{1}{2}\left[\frac{\tilde{\omega}_{2}\delta\omega_{2}}{\Delta_{m}^{2}}-\frac{\chi_{2}\delta\chi_{2}}{\Delta_{m}^{2}}-\frac{\chi_{1}\delta\chi_{1}}{\Delta_{m}^{1}}-\frac{1}{2}\left(\frac{\chi_{1}^{2}}{\tilde{\omega}_{1}}+\frac{\chi_{2}^{2}}{\tilde{\omega}_{2}}\right)\right]$ (41)
@@
$\sigma_{\rm{EP}}^{\epsilon}:=\sqrt{(8J_{m}r)^{2}+4\nu^{2}-4i\nu|\delta\gamma|+4\nu\mu-2i\mu|\delta\gamma|},$ (42)
with $\nu:=\frac{1}{2}\left(\frac{\chi_{2}^{2}}{\tilde{\omega}_{2}}-\frac{\chi_{1}^{2}}{\tilde{\omega}_{1}}\right)+\frac{\chi_{2}\delta\chi_{2}}{\Delta_{m}^{2}}-\frac{\tilde{\omega}_{2}\delta\omega_{2}}{\Delta_{m}^{2}}-\frac{\chi_{1}\delta\chi_{1}}{\Delta_{m}^{1}}$ and $\mu:=(2\delta\tilde{\omega}-{\rm i}\delta\Gamma_{\rm{eff}})$. In these expressions, $\delta\omega_{2}$ and $\delta\gamma$ are the two types of perturbation used in the main text. The former is induced by a mass deposition on the second mechanical resonator, while the latter comes from optical fluctuations or a mismatch damping of the two mechanical resonators. We stress that there can be also a perturbation induced by the fluctuation of the parametric driving which can be captured through $\delta\chi_{j}$ in Eq. 42. However, we have not considered such a perturbation in our investigation in the main text. We have also made some assumptions for a seek of simplicity, $\tilde{\omega}_{1}=\tilde{\omega}_{2}\equiv\omega_{m}$ and $\chi_{1}=\chi_{2}\equiv\chi$. With these approximations, we were able to derive different expressions for sensitivity and enhancement factor presented in the main text, which were qualitatively in good agreement with our results. In order to give a direct comparison between the two sensing scenarii, we plotted both their sensitivity and enhancement factor in Figure 5. The full line is from the sensing based on the shifting mechanism while the dashed line refers to splitting at the $\rm{EP}$. It can be clearly seen that the scheme based on the splitting performs better as discussed in the main text.
Figure 5: Sensing performances comparison. Sensitivities and the enhancement factors for the two sensing schemes when $\chi=0$. The used parameters are the same as those in Figure 1.
@@ Appendix C Transmission and output spectra
$\mathcal{O}(\omega)=\mathcal{D}(\omega)\left[{\rm i}\sqrt{\mathcal{K}}\mathcal{N}(\omega)+\sqrt{\gamma_{m}}\mathcal{O}^{\rm in}(\omega)\right],$ (43)
where $\mathcal{D}(\omega):=(-\mathcal{M}-{\rm i}\omega)^{-1}$. We are interested to obtain the spectra two optical modes at the output of the opto-mechanical cavity. According to the input-output theory, the operators for the output fields are related to the cavity and to the input noise operators by the relation $\delta\alpha^{\rm out}_{1}(\omega):=\sqrt{\kappa}\delta\alpha_{1}(\omega)-\delta\alpha^{\rm in}_{1}(\omega)$ and $\delta\alpha^{\rm out}_{2}(\omega):=\sqrt{\kappa}\delta\alpha_{2}(\omega)-\delta\alpha^{\rm in}_{2}(\omega)$. Then by using Eq.(A) and Eq.(43), the expressions for the two transmitted fields in Fourier domain are given by,
$\delta\alpha^{\rm out}_{1}(\omega)$ $=S_{11}(\omega)\delta\alpha^{\rm in}_{1}(\omega)+S_{12}(\omega)\delta\alpha^{in^{\dagger}}_{2}(\omega)+S_{13}(\omega)\delta\beta^{\rm in}_{1}(\omega)$
$+S_{14}(\omega)\beta^{\rm in}_{2}(\omega),$ (44a)
$\alpha^{out^{\dagger}}_{2}(\omega)$ $=S_{21}(\omega)\delta\alpha^{\rm in}_{1}(\omega)+S_{22}(\omega)\delta\alpha^{in^{\dagger}}_{2}(\omega)+S_{23}(\omega)\delta\beta^{\rm in}_{1}(\omega)$
$+S_{24}(\omega)\beta^{\rm in}_{2}(\omega)$ (44b)
@@
$S_{11}(\omega):=1-\dfrac{2\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},\,\,S_{12}(\omega):=-\dfrac{2{\rm i}\tilde{G}_{1}\tilde{J}_{m}\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},$ (45a)
$S_{13}(\omega):=\dfrac{2{\rm i}\tilde{G}_{1}\xi_{2}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},\,\,S_{14}(\omega):=\dfrac{2\tilde{G}_{1}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},$ (45b)
$S_{21}(\omega)=\dfrac{2{\rm i}\tilde{G}_{2}\tilde{J}_{m}\sqrt{\Gamma_{1}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},\,\,S_{22}(\omega):=1+\dfrac{2\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\Gamma_{2}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},$ (45c)
$S_{23}(\omega):=\dfrac{2\tilde{G}_{2}\tilde{J}_{m}\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},\,\,S_{24}(\omega):=-\dfrac{2{\rm i}\tilde{G}_{2}\xi_{1}(\omega)\sqrt{\gamma_{m}}}{\sqrt{\kappa}\xi_{\rm eff}(\omega)},$ (45d)
where $\xi_{\rm eff}(\omega)=\xi_{1}(\omega)\xi_{2}(\omega)+\tilde{J}^{2}_{m}$ with $\xi_{1}(\omega)=\Gamma^{1}_{\rm eff}/2+{\rm i}(\Delta^{1}_{m}-\omega)$ and $\xi_{2}(\omega)=\Gamma^{2}_{\rm eff}/2+{\rm i}(\Delta^{2}_{m}-\omega)$. The input quantum noise operators cavity and mechanical modes are characterized by the following correlations
$\langle\delta\alpha^{\rm in}_{j}(t),\delta\alpha^{in^{\dagger}}_{j}(t^{\prime})\rangle$ $=\delta(t+t^{\prime}),$ (46a)
$\langle\delta\beta^{in^{\dagger}}_{j}(t^{\prime}),\delta\beta^{\rm in}_{j}(t)\rangle$ $=N_{s}\delta(t+t^{\prime}),$ (46b)
$\langle\delta\beta^{\rm in}_{j}(t),\delta\beta^{\rm in}_{j}(t^{\prime})\rangle$ $=M_{s}\delta(t+t^{\prime}),\quad{\rm for}\,\,\,\,j\in\{1,2\},$ (46c)
where $N^{j}_{s}=(n^{j}_{m}+1)\sinh^{2}(r_{j})+n^{j}_{m}\cosh^{2}(r)$ and $M^{j}_{s}=(2n^{j}_{m}+1)\sinh(r_{j})\cosh(r_{j})$ with $n^{j}_{m}=[{\rm e}^{\hbar\omega_{j}/K_{B}T}-1]^{-1}$ is the mean thermal occupation number of the mechanical modes. Here we assumed that both mechanical resonators are kept at same temperature $T$. By using the expressions from Appendix C and Appendix C, one can derive the output spectra defined as
@@
$+\left(2N_{s_{2}}+1\right)\left|S_{24}(\omega)\right|{}^{2},$ (47b)
where $N_{s_{1,2}}:=(n_{m_{1,2}}+1)\sinh^{2}(r_{1,2})+n_{m_{1,2}}\cosh^{2}(r_{1,2})$.